OpenStudy (anonymous):

At the beginning of an experiment, a scientist has 268 grams of radioactive goo. After 90 minutes, her sample has decayed to 4.1875 grams. What is the half-life of the goo in minutes?______ Find a formula for G(t), the amount of goo remaining at time t. G(t)=______ How many grams of goo will remain after 65 minutes?_____

OpenStudy (misty1212):

HI again

OpenStudy (misty1212):

want to do this the easy way?

OpenStudy (anonymous):

yass plz

OpenStudy (misty1212):

or do you have to do this the \[A_0e^{-kt}\] way?

OpenStudy (misty1212):

either way we can do it

OpenStudy (anonymous):

thats basically Pert right?

OpenStudy (misty1212):

yeah

OpenStudy (anonymous):

then yeah use Pert

OpenStudy (misty1212):

\[P=268\]

OpenStudy (misty1212):

to find \(r\) as you called it, set \[4.1875=268e^{90r}\] and solve for \(r\)

OpenStudy (anonymous):

ok i know how to do that

OpenStudy (misty1212):

you get \[\frac{4.1875}{268}=e^{90r}\\ \ln(\frac{4.1875}{268}=90r\\ \frac{\ln(\frac{4.1875}{268})}{90})=r\]

OpenStudy (misty1212):

i get \[r=-0.0462098...\]

OpenStudy (anonymous):

same

OpenStudy (anonymous):

ok so now the formula is G(t)= 268e^-.04521(t)

OpenStudy (misty1212):

right and to find the half life set \[e^{-0.0462t}=\frac{1}{2}\] and solve for \(t\)

OpenStudy (anonymous):

ok thnx i got it now